OpenStudy (anonymous):

Solve using the square root method l. Round to nearest tenth if necessary. x^2 + 15 = 0

jimthompson5910 (jim_thompson5910):

Do you know about complex or imaginary numbers?

jimthompson5910 (jim_thompson5910):

In other words, have you seen \[\Large i=\sqrt{-1}\] before?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

ok so there will be no real solutions because \[\Large x^2 + 15 = 0\] \[\Large x^2 +15-15= 0-15\] \[\Large x^2 = -15\] \[\Large x = \sqrt{-15}\] Since you cannot take the square root of a negative number, there are no real solutions

jimthompson5910 (jim_thompson5910):

and you leave it at that because you haven't learned about imaginary numbers yet

OpenStudy (anonymous):

oh ok thanks

jimthompson5910 (jim_thompson5910):

yw